OpenStudy (caitlyn):

Can someone tell me the first step to solve this? 1. What is the factored form of 4x^2 + 12x + 5?

OpenStudy (tkhunny):

1) Don't be afraid of it. You can't break it. Try something -- anything. 2) Recognize that it just isn't that different from the other dozens of these you have already factored. Follow the same procedure and consider the same things. 3) You will need the factors of 4 and the factors of 5 1*4 = 4 2*2 = 4 1*5 = 5 4) Find the combination of factors and addition that produces 12 5) Wonder why you thought this one was so difficult.